OpenStudy (anonymous):

Definite Integrals: **MEDAL AND FAN** evaluate the given integral or state that it cannot be evaluated for the given information.

OpenStudy (anonymous):

\[\int\limits_{a}^{c} f(x)dx\] so what is it that I'm supposed to be doing? because in my book it has \[\int\limits_{a}^{c} g(x) dx =25\] and i have no idea how it go that??

OpenStudy (anonymous):

@jdoe0001 any help?

OpenStudy (anonymous):

i have here that \[\int\limits_{a}^{b}f(x)dx =7\] \[\int\limits_{a}^{b}g(x)dx=12\] and \[\int\limits_{b}^{c}g(x)dx=13\] so what would \[\int\limits_{a}^{c}f(x) be?\]

OpenStudy (anonymous):

Can you restate the problem? I am confused as which integrals are for which problem.

OpenStudy (anonymous):

those first three is what is already given to us. the fourth is the actual problem. so if we have \[\int\limits_{a}^{c}g(x)dx\] it equals 25 because \[\int\limits_{a}^{b}g(x)dx = 12 and \int\limits_{b}^{c}g(x)dx = 13\] @arkgolucky

OpenStudy (anonymous):

and 12+13 =25, get what its asking now?

OpenStudy (anonymous):

Oh okay I get it. I think drawing a picture will help you understand this problem.